the standard potential of the silver \silver bromed electrod is 0.073V.calculate the solubility product of silver bromide?
1
Expert's answer
2012-11-12T10:58:17-0500
Ksp = exp(z·E·F/R·T)
Ksp = exp(1·0.073·96485 / 8.314472· 298.15) = e^2.841 = 17.13